OpenStudy (anonymous):

Simplify completely 25 w to the sixth power over 10 w to the third power divided by 30 w squared over 5 w 1 over 15 w to the fourth power 5 w squared over 12 12 over 5 w squared 15w^4

OpenStudy (anonymous):

To make this more neat and clear, the problem is setup as this.. Simplify completely \[\frac{ 25w^6 }{ 10w^3 } \div \frac{ 30w^2 }{ 5w }\] The answers are: \[\frac{ 1 }{ 15w^4 }\] \[\frac{ 5w^2 }{ 12 }\] \[\frac{ 12 }{ 5w^2 }\] \[15w^4\] Correct?

OpenStudy (anonymous):

yea that is correct

OpenStudy (anonymous):

@Dakotafox79